OpenStudy (anonymous):

−3t−4/p=6/s solve for P

OpenStudy (anonymous):

is that (-3t-4)/p?? then p is.. p = s(-3t-4)/6

OpenStudy (anonymous):

how do you do it?

OpenStudy (anonymous):

is -3t /1-4/p=6/s

OpenStudy (anonymous):

but how do you do it ..please explain...